OpenStudy (anonymous):

Scoring a hole-in-one is the greatest shot a golfer can make. Once 5 professional golfers each made holes-in-one on the 7th hole at the same golf course at the same tournament. It has been found that the estimated probability of making a hole-in-one is (1/2494) for male professionals. Suppose that a sample of 5 professional male golfers is randomly selected. (a) What is the probability that at least one of these golfers makes a hole-in-one on the 10th hole at the same tournament? (b) What is the probability that all of these golfers make a hole-in-one on the 10th hole at the same tournament?

OpenStudy (anonymous):

at least one means not none. find the probability that none make it and subtract the result from one probability one makes it is 1/2494. probability that he doesn't make it is 2493/2494 probability that all 5 don't make it is \((\frac{2493}{2494})^5\)

OpenStudy (anonymous):

and so the answer you want is \(1-(\frac{2493}{2494})^5\) whatever that is

OpenStudy (anonymous):

probability they all do is \((\frac{1}{2494})^5\) which is very close to zero

OpenStudy (anonymous):

Yep! The second one is actually a zero. Thank you so much!

OpenStudy (anonymous):

Can you by any chance also please help me with this? Consider two people being randomly selected. (For simplicity, ignore leap years.) (a) What is the probability that two people were born on Monday? The method seems simple but I can't seem to get the right answer